Good(X): 1,2,3,4,5 (increasing)
Good(Y): 12,8,5,3,2(decreasing)

"Marginal rate of substitution (MRS) of X for Y will go on diminishing." I have a question on this sentence that is according to Good(X) and Good(Y) it is clearly seen that Y is decreasing or diminishing and X is increasing but the sentence tells 'MRS of X will go on diminishing' what is this? Could you please tell me on what basis and how can we say X is diminishing?
1
Expert's answer
2018-12-17T06:53:10-0500

Marginal rate of substitution (MRS) of X for Y is calculated using the formula

MRS=-∆y/∆x



In this case MRS is really diminishing, because

〖MRS〗_1=-((8-12))/((2-1) )=4


〖MRS〗_2=-((5-8))/((3-2) )=3


〖MRS〗_3=-((3-5))/((4-3) )=2


〖MRS〗_4=-((2-3))/((5-4) )=1